An airplane flying 60 m/s is accelerated uniformly at the rate of 0.5 . What is the velocity at the end 10 s
1
Expert's answer
2020-05-20T09:32:00-0400

Solution.

"v_0=60m\/s;"

"a=0.5 m\/s^2;"

"t=10s;"

"v-?;"

Acceleration at uniformly accelerated motion is determined by the formula:

"a=\dfrac{v-v_0}{t};\implies v=v_0+at;"

"v=60m/s+0.5m/s^2\sdot10s=65m/s;"

Answer: "v=65m\/s."